# Quantile Activation: departing from single point estimation for better generalization across distortions

### v1

@@ 3 Training with QAct
One can use any differentiable loss function to train with quantile activation. We specifically experiment with the standard Cross-Entropy Loss, Triplet Loss, and the recently proposed Watershed Loss [6] (see section 4). However, if one requires that the boundaries between classes adapt to the distribution, then learning similarities instead of boundaries can be beneficial. Both Triplet Loss and Watershed Loss fall into this category. We see that learning similarities does have slight benefits when considering the embedding quality.
@@
Observe that the proposed QAct (without normalization) returns the values in $[0,1]$ which can be interpreted as probabilities. Hence, one can also use this for the classification layer. Nonetheless, two changes are required – (i) Traditional softmax used in conjunction with negative-log-likelihoood loss already considers “relative” activations of the classification in normalization. However, QAct does not. Hence, one should use Binary-Cross-Entropy loss with QAct, which amounts to one-vs-rest classification. (ii) Also, unlike a neuron in the middle layers, the bias of the neuron in the classification layer depends on the class imbalance. For instance, with $10$ classes, one would have only $1/10$ of the samples labelled $1$ and $9/10$ of the samples labelled 00. To address this, we require that the median of the outputs be at $0.9$, and hence weight the positive class with $0.9$ and the negative class with $0.1$ respectively. In this article, whenever QAct is used, we use this approach for inference.
We observe that (figures 13 and 14) using quantile classifier on the learned features in general improves the consistency of the calibration error and also leads to the reducing the calibration error. In this article, for all networks trained with quantile activation, we use quantile classifier to compute the accracies/calibration errors.
@@ 4 Evaluation
To summarize, we make the following changes to the existing classification pipeline – (i) Replace the usual ReLU activation with QAct and (ii) Use triplet or watershed loss instead of standard cross entropy loss. We expect this framework to learn context dependent features, and hence be robust to distortions. (iii) Also, use quantile classifier to train the classifier on the embedding for better calibrated probabilities.
@@
To evaluate our approach, we consider the two datasets developed for this purpose – CIFAR10C, CIFAR100C, TinyImagenetC [15], MNISTC[25]. These datasets have a set of 15 distortions at 5 severity levels. To ensure diversity we evaluate our method on 4 architectures – (overparametrized) LeNet, ResNet18[14] (11M parameters), VGG[35](15M parameters) and DenseNet [18](1M parameters). The code to reproduce the results can be found at https://anonymous.4open.science/r/QuantAct-2B41.
@@
To our knowledge, there exists no other framework which proposed classification based on context distribution. So, for comparison, we consider standard ReLU activation [10], pReLU [13], and SELU [20] for all the architectures stated above. Also, we compare our results with DINOv2 (small) [27] (22M parameters) which is current state-of-the-art for domain generalization. Note that for DINOv2, architecture and datasets used for training are substantially different (and substantially larger) from what we consider in this article. Nevertheless, we include the results for understanding where our proposed approach lies on the spectrum. We consider the small version of DINOv2 to match the number of parameters with the compared models.
@@
We consider four metrics – Accuracy (ACC), calibration error (ECE) [23] (both marginal and Top-Label) and mean average precision at $K$ (MAP@K) to evaluate the embedding. For the case of ReLU/pReLU/SELU activation with Cross-Entropy, we use the logistic regression trained on the train set embeddings, and for QAct we use the calibrated linear classifier, as proposed above. We do not perform any additional calibration and use the probabilities. We discuss a selected set of results in the main article. Please see appendix C for more comprehensive results.
Calibration error measures the reliability of predicted probabilities. In simple words, if one predicts 100 samples with (say) probability 0.7, then we expect 70 of the samples to belong to class $1$ and the rest to class 00. This is measured using either the marginal or top-label calibration error. We refer the reader to [23] for details, which also provides an implementation to estimate the calibration error.
@@
To verify that most existing activations do not share the robustness property of QAct, we compare QAct with other activations in figures 4(b) and 4(c). We observe that QAct is greatly more robust with respect to distortions in both accuracy and calibration error than other activation functions.
@@ 5 Conclusion And Future Work
    The key idea in our proposed approach is that the quantiles capture the distribution of each neuron from the batch of samples, providing outputs accordingly. This poses a challenge for inference, and we have discussed two potential solutions: (i) remember the quantiles and density estimates for single sample evaluation, or (ii) ensure that a batch of samples from the same distribution is processed together. We adopt the latter method in this article. An alternative approach would be to learn the distribution of each neuron using auxiliary loss functions, adjusting these distributions to fit the domain at test time. This gives us more control over the network at test time compared to current workflows.
@@
    On the theoretical side, there is an interesting analogy between our quantile activation and how a biological neuron behaves. It is known that when the inputs to a biological neuron change, the neuron adapts to these changes [7]. Quantile activation does something very similar, which leads to an open question – can we establish a formal link between the adaptability of a biological neuron and the accuracy of classification systems?
@@
    Another theoretical direction to explore involves considering distributions not just at the neuron level, but at the layer level, introducing a high-dimensional aspect to the problem. The main challenge here is defining and utilizing high dimensional quantiles, which remains an open question [21].
Broad Impact:
In this article, we propose an approach to maintain calibration and generalization across small distortions. While, we do not foresee any direct societal consequences of our work, we expect the potential future consequences of the technique to reduce the bias in the following ways – (i) Since we do not assume normal distribution, our approach is likely to handle long tails better than existing methods. This would help in reducing the dataset bias where marginal groups are less represented. (ii) Note that the output of each QAct layer is a uniform distribution. This can allow us to understand the working of each layer in isolation and possibly reduce the black-box nature of the current classification systems. (iii) Moreover, by directly modifying the context distribution of each neuron, one can easily make the networks more reliable without resorting to expensive re-training the entire network.
@@ Appendix C Extended Results Section
Figure 6 shows the corresponding results. The first experiment compares QAct + watershed with ReLU + Cross-Entropy on two standard networks – ResNet18 and DenseNet. With respect to accuracy, we observe that while at severity 00, ReLU + Cross-Entropy slightly outperforms QAct + watershed, as severity increases QAct + watershed is far more stable. We even outperform DinoV2(small) (22M parameters) at severity $5$. Moreover, with respect to calibration error, we see a consistent trend across distortions. As [5] argues, this helps in building more robust systems compared to one where calibration error increases across distortions.
@@
Figures 10 and 11 shows the comparison of QAct with ReLU, pReLU and SeLU. We observe the same trend across ReLU, pReLU and SeLU, while QAct is far more stable across distortions.
@@
Figures 13 and 14 shows the effect of quantile classifier on standard ResNet10/DinoV2 outputs with CIFAR10C/CIFAR100C datasets. While the accuracy values are almost equivalent, we observe a “flatter” trend of the calibration errors, sometimes reducing the error as in the case of CIFAR100C.
@@ Appendix D Watershed Loss
The authors in [6] proposed a novel classifier – watershed classifier, which works by learning similarities instead of the boundaries. Below we give the brief idea of the loss function, and refer the reader to the original paper for further details.
@@
    Watershed $\text{Watershed Loss}=\sum_{i=1}^{n_{\text{samples}}}\sum_{l=1}^{L}I[y_{i}=l]\log\left(\frac{\exp\left(-\|f_{\theta}({\bm{x}}_{i})-f_{\theta}({\bm{x}}_{i,l,1nn})\|\right)}{\sum_{j=1}^{L}\exp\left(-\|f_{\theta}({\bm{x}}_{i})-f_{\theta}({\bm{x}}_{i,j,1nn})\|\right)}\right)$ (10)
@@
If one uses (say) Cross Entropy loss, then the features learned would be such that the classes are linearly separable. Contrast this with watershed, which instead learns a similarity between two points in a translation invariant manner.
@@
Observe that the watershed loss is very similar to metric learning losses. The authors in [6] claim that this offers better generalization, and show that this is consistent with 1NN classifier. Moreover, they show that this classifier (without considering $f_{\theta}$) has a VC dimension which is equal to the number of classes. While metric learning losses are similar, there is no such guarantee with respect to classification. This motivated our choice of using watershed loss over other metric learning losses.

### v2

@@ 3 Training with QAct
One can use any differentiable loss function to train with quantile activation. We specifically experiment with the standard Cross-Entropy Loss, Triplet Loss, and the recently proposed Watershed Loss in (Challa et al., 2024) (see section 4). However, if one requires that the boundaries between classes adapt to the distribution, then learning similarities instead of boundaries can be beneficial. Both Triplet Loss and Watershed Loss fall into this category. We see that learning similarities does have slight benefits when considering the embedding quality.
@@
Observe that the proposed QAct (without normalization) returns the values in $[0,1]$ which can be interpreted as probabilities. Hence, one can also use this for the classification layer. Nonetheless, two changes are required – (i) Traditional softmax used in conjunction with negative-log-likelihood loss already considers “relative” activations of the classification in normalization. However, QAct does not. Hence, one should use Binary-Cross-Entropy loss with QAct, which amounts to one-vs-rest classification. (ii) Also, unlike a neuron in the middle layers, the bias of the neuron in the classification layer depends on the class imbalance. For instance, with $10$ classes, one would have only $1/10$ of the samples labelled $1$ and $9/10$ of the samples labelled 00. To address this, we require that the median of the outputs be at $0.9$, and hence weight the positive class with $0.9$ and the negative class with $0.1$ respectively. In this article, whenever QAct is used, we use this approach for inference.
We observe that (figures 14 and 15) using quantile classifier on the learned features in general improves the consistency of the calibration error and also leads to the reducing the calibration error. In this article, for all networks trained with quantile activation, we use quantile classifier to compute the accuracy/calibration errors.
@@ 4 Evaluation
To summarize, we make the following changes to the existing classification pipeline – (i) Replace the usual ReLU activation with QAct and (ii) Use triplet or watershed loss instead of standard cross-entropy loss. We expect this framework to learn context dependent features, and hence be robust to distortions. (iii) Also, use quantile classifier to train the classifier on the embedding for better calibrated probabilities.
@@
To evaluate our approach, we consider the datasets developed for this purpose – CIFAR10C, CIFAR100C, TinyImagenetC (Hendrycks & Dietterich, 2019), MNISTC (Mu & Gilmer, 2019). These datasets have a set of 15 distortions at 5 severity levels. To ensure diversity we evaluate our method on 4 architectures – (overparametrized) LeNet, ResNet18 (He et al., 2016) (11M parameters), VGG (Simonyan & Zisserman, 2015)(15M parameters) and DenseNet (Huang et al., 2017) (1M parameters). The code to reproduce the results can be found at https://anonymous.4open.science/r/QuantAct-534C.
@@
To our knowledge, there exists no other framework which proposes classification based on context distribution. So, for comparison, we consider standard ReLU activation (Fukushima, 1970), pReLU (He et al., 2015), and SELU (Klambauer et al., 2017) for all the architectures stated above. Also, we compare our results with DINOv2 (small) (Oquab et al., 2023) (22M parameters) which is current state-of-the-art for domain generalization. Note that for DINOv2, architecture and datasets used for training are substantially different (and substantially larger) from what we consider in this article. Nevertheless, we include the results for understanding where our proposed approach lies on the spectrum. We consider the small version of DINOv2 to match the number of parameters with the compared models.
@@
We consider four metrics – Accuracy (ACC), calibration error (ECE) (Kumar et al., 2019) (both marginal and Top-Label) and mean average precision at $K$ (MAP@K) to evaluate the embedding. For the case of ReLU/pReLU/SELU activation with Cross-Entropy, we use the logistic regression trained on the train set embeddings, and for QAct we use the calibrated linear classifier, as proposed above. We do not perform any additional calibration and use the probabilities. We discuss a selected set of results in the main article. Please see appendix C for more comprehensive results.
Calibration error measures the reliability of predicted probabilities. In simple words, if one predicts 100 samples with (say) probability 0.7, then we expect 70 of the samples to belong to class $1$ and the rest to class 00. This is measured using either the marginal or top-label calibration error. We refer the reader to (Kumar et al., 2019) for details, which also provides an implementation to estimate the calibration error.
@@
To verify that most existing activations do not share the robustness property of QAct, we compare QAct with other activations in figures 5(b) and 5(c). We observe that QAct is greatly more robust with respect to distortions in both accuracy and calibration error than other activation functions.
@@ 5 Conclusion And Future Work
    The key idea in our proposed approach is that the quantiles capture the distribution of each neuron from the batch of samples, providing outputs accordingly. This poses a challenge for large datasets, and we have discussed two potential solutions: (i) remember the quantiles and density estimates for single sample evaluation, or (ii) ensure that a batch of samples from the same distribution is processed together. We adopt the latter method in this article. An alternative approach would be to learn the distribution of each neuron using auxiliary loss functions, adjusting these distributions to fit the domain at test time. This gives us more control over the network at test time compared to current workflows. If the networks are very large, where batch sizes cannot be big – there exists several strategies such as checkpointing to implicitly increase the batch size.
@@
    On the theoretical side, there is an interesting analogy between our quantile activation and how a biological neuron behaves. It is known that when the inputs to a biological neuron change, the neuron adapts to these changes (Clifford et al., 2007). Quantile activation does something very similar, which leads to an open question – can we establish a formal link between the adaptability of a biological neuron and the accuracy of classification systems?
@@
    Another theoretical direction to explore involves considering distributions not just at the neuron level, but at the layer level, introducing a high-dimensional aspect to the problem. The main challenge here is defining and utilizing high dimensional quantiles, which remains an open question (Koenker, 2005).
@@ Appendix C Extended Results Section
Figure 7 shows the corresponding results. The first experiment compares QAct + watershed with ReLU + Cross-Entropy on two standard networks – ResNet18 and DenseNet. With respect to accuracy, we observe that while at severity 00, ReLU + Cross-Entropy slightly outperforms QAct + watershed, as severity increases QAct + watershed is far more stable. We even outperform DINOv2(small) (22M parameters) at severity $5$. Moreover, with respect to calibration error, we see a consistent trend across distortions. As (Challa et al., 2023) argues, this helps in building more robust systems compared to one where calibration error increases across distortions.
@@
Figures 11 and 12 shows the comparison of QAct with ReLU, pReLU and SeLU. We observe the same trend across ReLU, pReLU and SeLU, while QAct is far more stable across distortions.
@@
Figures 14 and 15 shows the effect of quantile classifier on standard ResNet10/DinoV2 outputs with CIFAR10C/CIFAR100C datasets. While the accuracy values are almost equivalent, we observe a “flatter” trend of the calibration errors, sometimes reducing the error as in the case of CIFAR100C.
Measuring Robustness of QAct:
To measure the robustness of the proposed method we use the metric Acc@Dist_i - Acc@Dist_j which measures the drop in accuracy when the distortion severity is increased from $i\to j$. A method is considered to be better if the values are lower, i.e if the drop in accuracy is smaller than comparitive methods. Tables 1 and 2 shows the comparison between ReLU, QAct and DinoV2(s). We see that, in all the cases QAct outperforms both ReLU and DinoV2(s) at all possible $i\to j$.
0$\to$1 0$\to$2 0$\to$3 0$\to$4 0$\to$5 1$\to$2 1$\to$3 1$\to$4 1$\to$5
Dataset Model/Method
CIFAR10 Resnet18-ReLU 6.70 11.57 17.27 24.82 35.49 4.87 10.57 18.12 28.79
DinoV2(s) 6.92 10.66 15.74 21.58 30.93 3.74 8.82 14.67 24.02
Resnet18-Quant 4.42 6.88 9.22 12.42 16.86 2.46 4.80 8.00 12.44
CIFAR100 Resnet18-ReLU 10.25 17.85 23.78 29.81 36.87 7.59 13.52 19.56 26.62
DinoV2(s) 10.64 16.39 22.54 28.71 38.02 5.74 11.89 18.06 27.37
Resnet18-Quant 5.83 9.07 12.20 15.93 21.00 3.24 6.38 10.10 15.17
TinyImagenet Resnet18-ReLU 5.51 9.58 14.90 19.13 21.49 4.07 9.39 13.62 15.97
DinoV2(s) 9.39 14.73 23.30 33.78 42.52 5.34 13.91 24.39 33.13
Resnet18-Quant 3.27 4.86 7.26 10.43 13.18 1.59 3.99 7.16 9.90
Table 1: Measuring the drop in accuracy
2$\to$3 2$\to$4 2$\to$5 3$\to$4 3$\to$5 4$\to$5
Dataset Model/Method
CIFAR10 Resnet18-ReLU 5.70 13.25 23.92 7.55 18.22 10.67
DinoV2(s) 5.08 10.93 20.28 5.85 15.20 9.35
Resnet18-Quant 2.34 5.54 9.98 3.20 7.64 4.45
CIFAR100 Resnet18-ReLU 5.93 11.97 19.03 6.04 13.10 7.06
DinoV2(s) 6.15 12.32 21.63 6.17 15.48 9.31
Resnet18-Quant 3.13 6.86 11.92 3.73 8.79 5.06
TinyImagenet Resnet18-ReLU 5.32 9.55 11.90 4.23 6.58 2.35
DinoV2(s) 8.57 19.05 27.79 10.48 19.22 8.74
Resnet18-Quant 2.40 5.56 8.31 3.17 5.92 2.75
Table 2: Measuring the drop in accuracy (contd. from table 1)
@@ Appendix D Watershed Loss
The authors in (Challa et al., 2024) proposed a novel classifier – watershed classifier, which works by learning similarities instead of the boundaries. Below we give the brief idea of the loss function, and refer the reader to the original paper for further details.
@@
    Watershed $\text{Watershed Loss}=\frac{-1}{n_{\text{samples}}}\sum_{i=1}^{n_{\text{samples}}}\sum_{l=1}^{L}I[y_{i}=l]\log\left(\frac{\exp\left(-\|f_{\theta}({\bm{x}}_{i})-f_{\theta}({\bm{x}}_{i,l,1nn})\|\right)}{\sum_{j=1}^{L}\exp\left(-\|f_{\theta}({\bm{x}}_{i})-f_{\theta}({\bm{x}}_{i,j,1nn})\|\right)}\right)$ (11)
@@
If one uses (say) cross-entropy loss, then the features learned would be such that the classes are linearly separable. Contrast this with watershed, which instead learns a similarity between two points in a translation invariant manner.
@@
Observe that the watershed loss is very similar to metric learning losses. The authors in (Challa et al., 2024) claim that this offers better generalization, and show that this is consistent with 1NN classifier. Moreover, they show that this classifier (without considering $f_{\theta}$) has a VC dimension which is equal to the number of classes. While metric learning losses are similar, there is no such guarantee with respect to classification. This motivated our choice of using watershed loss over other metric learning losses.